Many years ago, Julie deposited $14,490 into a savings account. Today, the balance of the account is $62,625. How long ago diMultiple Choice 29.51 years 0 29.72 years 29.72 years 30.13 years 30.00 years 29.67 years
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

62625=14490*(1.05)^n

(62625/14490)=(1.05)^n

Taking log on both sides;

log (62625/14490)=n*log 1.05

n=log (62625/14490)/log 1.05

=30 years(Approx).